Question Completion Status: QUESTION 20 7.00000 points (Extra Credit) Save Ane GS construction was awarded a contract to build a commercial building in 2008. The contract price was $4,800,000 for four years. Please complete the table based on the following transactions. 4,800,000.00 $ 4,500,000.00 Contract Price Estimated cost Start Date Complete Date Balance sheet date une, 06 October, 2011 31-Dec Costs to date Estimated costs to complete Progress billing during year Cash collected during year 650,000.00 s 2.160,000.00 3.3700000 4490.000.00 $ 2,340,000.00 5 3,850,000.00 2,100,000.00 S 1.650,000.00 $650,000.00 S 800,000.00 1,600,000.00 What is Revenue of 2008 in Percentage of Completion method? 8 9 alt ctrl
media%2Fd2a%2Fd2aa920c-8774-47b6-9bd5-ca
Answer #1

percent completed = (cost to date) / total estimation

= (650,000 / 4500,000 )*100 = 14.44%

Revenue = (completed % ) * contract price = (14.44 / 100) (4800000) = 693,333

Hence Revenue in year 2008 is $693,333